OpenStudy (anonymous):

tan (1/2)(sin^-1(2x/(1+x^2))+cos^-1((1-y^2)/(1+y^2)))

OpenStudy (anonymous):

\[\tan \frac{1}{2}\sin^-1(\frac{2x}{1+x^2})+\cos^-1(\frac{1-y^2}{1+y^2})\]

OpenStudy (anonymous):

sorry! correction below :

OpenStudy (anonymous):

\[\tan (1/2)(\sin^-1\frac{2x}{1+x^2}+\cos^-1\frac{1-y^2}{1+y^2})\]